OpenStudy (anonymous):

2x+y=-8 x-2y=-6 solve the system of linear equations

OpenStudy (anonymous):

x=-8-y/2 from eq1 put value of x in eq2 you'll get x .. then put actual value of x in any of the 2 eq you'll get y

OpenStudy (anonymous):

or second way is to solve them by matrices using echelon form .

OpenStudy (anonymous):

this is wut i got y=45 and x=-22/5

OpenStudy (anonymous):

answers are wrong they ain't satisfying the eq

OpenStudy (lalaly):

2x+y=-8.........(1) x-2y=-6 .........(2) multiply equation 1 by 2 4x+2y=-16 x-2y =-6 add both eqns 5x=-10 x=-2 substitute and find y

OpenStudy (lalaly):

oops mistake adding both equans 5x=-22

OpenStudy (lalaly):

so u are right @Astemmene

OpenStudy (anonymous):

lalaly solutions aint satisfying the eqs

OpenStudy (anonymous):

yeah mee!!! :D

OpenStudy (lalaly):

let me check

OpenStudy (lalaly):

the y should = 4/5 not 45

OpenStudy (anonymous):

woops i wrote it out wrong but thats wut i ment :P

OpenStudy (lalaly):

lol its ok xD

OpenStudy (anonymous):

i m getting x= -14/3 y = 4/3